Errata to

Abstract Evolution Equations, Periodic Problems and Applications

by D. Daners and P. Koch Medina April 28, 1999

p.21 l.4 smooth Jordan curve should read piecewise smooth Jordan curve

p.30 l.11 ẇ_{t}(x) = \frac{1}{t}(−\frac{n}{2}+\frac{|x|^{2}}{4t} ) should read ẇ_{t}(x) = \frac{1}{t}(−\frac{n}{2}+\frac{|x|^{2}}{4t} )w_{t}(x)

p.40 l.14 an isomorphismus, if T ∈ Isom(E_{0}, F_{0}). should read an isomor-

phismus if T ∈ Isom(E_{0}, F_{0}) and T ∈ L(E_{1}, F_{1}) is surjective.

p.41 l.23 by (F 2) it holds that… should read by (I2) it holds that… p.49 l.6 max{∥f∥_{BC(S_{0},E_{0})}∥, ∥f∥_{BC(S_{1},E_{1})}} should read

max{∥f∥_{BC(∂_{0}S,E_{0})}∥, ∥f∥_{BC(∂_{1}S,E_{1})}}

p.105 l.19 By (10.5) and (10.3) … should read By (10.5) and (10.6) … p.105 l.20,21 In formula (10.39) e^{−(t−τ)} should read e^{−n(t−τ)} p.113 l.28 ∥U(t, s)∥α, α should read ∥U(t, s)∥_{α,α} p.125 l.21 φ(t) = e^{μ(λ)t}φ_{0} should read φ(t) = e^{μ(λ)t}U(t, 0)φ_{0} p.164 l.5 P hi(u, λ) should read Φ(u, λ) p.170 l.3 the second term on the right hand side of (18.12) should read

∫_{s}^{t} U(t, τ)D_{3}g(τ, u(τ; s, x, λ), λ) dτ

p.203 l.3 +k(t)Δ should read −k(t)Δ p.203 l.19 … X_{0}-realization of k(t)Δ … should read … X_{0}-realization of

−k(t)Δ …

p.204 l.10 +k(t)Δ should read −k(t)Δ p.205 l.11 +k(t)Δ should read −k(t)Δ p.206 l.3 … supersolutions of (24.1) … should read … supersolutions of

(25.1) …

p.206 l.8 +k(t)Δ should read −k(t)Δ p.206 l.10 a, b ∈ BUC^{μ, \frac{μ}{2}} (ℝ^{N}×[0, T]) … should read m, b ∈ BUC^{μ, \frac{μ}{2}} (ℝ^{N}×

[0, T]) …

p.206 l.21 +k(t)Δ should read −k(t)Δ p.207 l.3 a(x, t) ≤ −γ′ < 0 should read m(x, t) ≤ −γ′ < 0 p.207 l.6 +k(t)Δ should read −k(t)Δ p.207 l.20,25,26 (25.7) should read (25.6) p.208 l.2 On two occasions (25.7) should read (25.6) p.208 l.4 (25.8) should read (25.9) and +k(t)Δ should read −k(t)Δ p.208 l.8 zero solution of (25.8) should read zero solution of (25.9) p.226 l.9 On two occasions lim_{r↘0} sup_{\begin{matrix}x=neqy \\ |x−y|≤r \end{matrix}} should read lim_{r↘0} sup_{\begin{matrix}x≠y \\ |x−y|≤r \end{matrix}}

p.239 l.3 W^{2,1}(Ω × [0, T]) should read W_{p}^{2,1}(Ω × [0, T]) p.247 l.25 J.L. Lions should read P.L. Lions p.248 l.11,13 Reidlinger should read Redlinger